OpenStudy (anonymous):

solve the equation: y = log(3 + x )− log( x − 5)=log3

OpenStudy (a_clan):

log m - log n = log(m/n)

OpenStudy (anonymous):

log[(3+x)/(x-5)]=log3 then: (3+x)/(x-5) = 3 can you solve it from this point?

OpenStudy (anonymous):

3+x=3x-15 2x=18 x=9